OpenStudy (anonymous):

23.621 to 29.9% of what number

OpenStudy (amistre64):

part/percent = total

OpenStudy (anonymous):

23.621/x=.299

OpenStudy (amistre64):

calculator says 79

OpenStudy (anonymous):

=79

OpenStudy (anonymous):

\[\frac{23.621}{x}\times 100=29.9\] so \[\frac{2362.1}{29.9}=x=79\]